OpenStudy (anonymous):

Choose the correct first step(s) to solve the system using elimination. 4x + 7y = 6 and 6x + y = 2 (Points : 1) 4x + y = 6 → 4x + y = 6 (6x + y = 2)(7)→ 42x +7y = 14 A) 4x + 7y = 6x + y B) y = -6x + 20 → 4x + 7(-6x + 20) = 6 C) (4x + 7y) + (6x + y) = 8 D) (4x + 7y) - (6x + y) = 4

OpenStudy (anonymous):

last option. .

OpenStudy (anonymous):

elimination is substitution. .

OpenStudy (anonymous):

answer is D)

OpenStudy (anonymous):

4x + 7y = 6 EQ(1) and 6x + y = 2 lets try solving this one D) (4x + 7y) - (6x + y) = 4 (4x + 7y) - (6x + y) = 4 (4x + 7y) - 6x - y = 4 -2x+6y=4 .. multiplying this by 2 we get -4x+12y=8 EQ(2) add this to EQ(1) 4x + 7y = 6 ------------ 19y=14 y=14/19 ans you can now subs this to get x